Solve the following systems of equation by substitution:

4. x=2y
3x-2y=24

Answer by mananth(16946)

4. x=2y
3x-2y=24


4x=2y
divide by 2
y=2x
substitute y in second equation
3x-2(2x)=24
3x-4x=24
-x=24
x=-24
plug x in first equation
4x=2y
4(-24)=2y
y=-48